# SkyAI Sim: An Open-Source Simulation of UAV Aerial Imaging from Satellite Data

### v1

@@ Abstract
Capturing real-world aerial images for vision-based navigation (VBN) is challenging due to limited availability and conditions that make it nearly impossible to access all desired images from any location. The complexity increases when multiple locations are involved. The state of the art solutions, such as flying a UAV (Unmanned Aerial Vehicle) to take pictures or using existing research databases, have significant limitations. SkyAI Sim offers a compelling alternative by simulating a UAV to capture bird’s-eye view satellite images at zero-yaw with real-world visible-band specifications. This open-source tool allows users to specify the bounding box (top-left and bottom-right) coordinates of any region on a map. Without the need to physically fly a drone, the virtual Python UAV performs a raster search to capture satellite images using the Google Maps Static API. Users can define parameters such as flight altitude, aspect ratio and diagonal field of view of the camera, and the overlap between consecutive images. SkyAI Sim’s capabilities range from capturing a few low-altitude images for basic applications to generating extensive datasets of entire cities for complex tasks like deep learning. This versatility makes SkyAI a valuable tool for not only VBN, but also other applications including environmental monitoring, construction, and city management. The open-source nature of the tool also allows for extending the raster search to other missions. A dataset of Memphis, TN has been provided along with this simulator, partially generated using SkyAI and, also includes data from a 3D world generation package for comparison.
Keywords Vision-Based Navigation  $\cdot$ UAV  $\cdot$ Satellite Image  $\cdot$ Aerial Imaging  $\cdot$ Simulation  $\cdot$ Dataset  $\cdot$ Google Map API
@@ 1 Introduction
Figure 1: Image captured at coordinates (35.128039, -89.799163) at above ground level altitude ($AGL$) of 126 m. Image captured (a) by a UAV and simulated (b-d) with a field of view ($FOV$) of 78.8 degrees and aspect ratio ($AR$) of 4:3. Figures b-d represent simulation using three different maptypes on Google Map Static API.
@@ 1.1 Application
Aerial imagery is essential for many applications, such as construction, city administration, environmental monitoring, and VBN. In artifitial intelligence (AI) and machine learning (ML), these images provide training data for models, aiding in accurate analysis and decision-making. In VBN, they help autonomous systems, like drones, identify obstacles and landmarks [24]. Environmental monitoring relies on aerial imagery to assess changes in land use, track deforestation, and monitor wildlife habitats [16]. In construction, they enhance project management by offering site overviews, improving resource allocation, and ensuring safety [17]. For city management, they facilitate urban planning, infrastructure maintenance, and disaster response [11]. Additionally, aerial images are key in AI/ML for developing object detection, classification, and segmentation algorithms [15, 18, 22].
@@ 1.2 Related Work
Acquiring real-world aerial images is challenging due to constraints. The limited availability of images is a primary issue, as it is often impractical to obtain the required aerial views from specific locations. This difficulty is compounded when there is a need to cover multiple locations on Earth. Deploying Unmanned Aerial Vehicles (UAVs), have inherent limitations. For instance, flying a UAV to capture aerial images involves logistical complexities, significant time investment, and financial costs. Moreover, UAV operations are subject to regulatory restrictions, such as airspace regulations and flight permissions, as well as environmental factors like weather conditions that can impede flight missions.
While there are many valuable datasets, they have limitations. Every dataset is designed for a specific applications such as segmentation, object detection, and more. Reference [12] gives a comprehensive collection of many of these datasets. There are real-world aerial images such as FloodNet [20, 21]. Floodnet includes 2343 samples of flooded areas and provides a comprehensive dataset for classification, semantic segmentation, and visual question answering. On the other hand, many datasets are derived from satellite data. The University1652-Baseline dataset [26, 25] provides both drone-based and satellite-view geo-localization datasets. While these datasets are highly valuable for their specific tasks, they are limited by the parameter selection of the authors. They cannot replace the benefit of choosing preferred locations and altitudes as a real drone mission would.
Although datasets are useful, they may not cover the exact scenarios needed for comprehensive testing or algorithm development. Simulators offer a controlled environment to generate large amounts of synthetic data, customize conditions, and adapt to new models without the constraints of physical data collection. They also provide repeatability, safety, and cost-effectiveness, addressing issues like incomplete data or challenging environments that might be difficult to replicate with real-world datasets.
SkyAI Sim  [^1] https://github.com/JacobsSensorLab/SkyAI-Sim addresses these challenges by offering a versatile tool to capture geo-tagged simulated aerial images. It emulates a UAV to capture bird’s-eye view satellite images that adhere to real-world visible-band specifications, with the drone facing north at a zero yaw as represented in Figure 1. A bird’s-eye view is a 90-degree tilt, means the drone’s camera is looking straight down, orthogonal to the ground. However, this limitation is not significant for most applications as many applications predominantly use bird’s-eye view images. With SkyAI, users can simulate various missions using the Google Maps Static API in Python, by specifying the coordinates of any region on a map. Furthermore, if the users have the tools to collect their own drone-view images, SkyAI expands the capabilities of the University1652 dataset to any desired location. This enables them to perform tasks based on both drone-view and satellite-view images.
1.3 Google Maps Static API
The Google Maps Static API allows embedding a static map image on a webpage without needing JavaScript or dynamic content loading. The API generates a map based on specified settings and provides it as an image through a standard HTTP request with URL parameters. This simplifies the integration of maps into webpages, offering customization options including map center, zoom level, size, and type, making it versatile for adding geographic visualizations without dynamic interactions [3]. The API uses the Web Mercator Projection, a cylindrical map projection that preserves angles and shapes but distorts size and distance, especially near the poles [5, 9]. Understanding this projection aids in accurately mapping dimensions from the API to the real world.
@@ 2.1.1 Single Image Capture
The process of capturing a map from the Google Maps Static API [3], involves understanding the relationship between the zoom level, image pixel size, and geolocation bounds. The API allows one to specify the zoom level of the map. The zoom level determines the scale of the map and, consequently, the resolution of the captured image. Higher zoom levels provide more detail, while lower zoom levels cover larger areas with less detail. As follows, this section explains how the simulation parameters $AGL,FOV,AR$ are converted to parameters required to capture satellite data including $z$ and $res_{x,y}$  [^2] The functions introduced in ”Single Image Capture” section endnotes are located in the ”src/utils” folder within the ”geo_helper” package..
@@
    The Google Maps JavaScript API [1] The Google Maps JavaScript API’s getBounds() function provides the latitude and longitude boundaries of the visible map area based on the central coordinates, zoom level, and map size in pixels ($res_{x,y}$, also referred to as the resolution of the map). To simulate a geographic area from a central point, zoom level, and map dimensions, we calculate the bounding box coordinates using Mercator projection principles. The following steps detail this calculation process. 555”calc_bbox_api” function implements bounding box calculation in Python.
@@
  • Get Static Map Image
    Finally, once the optimal zoom level is determined, the API is used to capture the image. This is done by setting the map’s center, zoom level, and image pixel size on the API’s URL. For realistic images, the map type is set to "satellite" by default. However, the program is not limited to capturing realistic images and different map types can be requested [4]. Using "satellite" data, the zoom level can range between 0 and 22 as mentioned in Table 1. It’s important to note that the zoom level limits vary across different map types. The directory for where to save the captured data has to be determined. A flowchart of the process is available in Figure 3. An API key should be generated to access the API  [^7] The ”init_static_map” takes in the converted information and returns a URL response in the output. It is important to note that this function removes all the street names and text labels from the downloaded image. This can be undone by removing the ”style” parameter in the function. API-key can be fed to the program directly, or specified in a file named ”hidden_file.py” in the source code directory. This key should not be shared publicly and should stay secure. ,  [^8] The URL output can be downloaded locally using ”get_static_map_image” function. The ”retry” parameter determines the number of retries if the download fails..
@@ 2.1.2 Raster Image Mission
A logical extension of a single image capture, can define various missions. The main purpose of SkyAI Sim is to capture raster images of a specified area. In this mission, the objective is to complete a zero-tilt raster image capture, as shown in Fig. 2. The capture begins once the bounding box coordinates of the mission’s rectangular region (the map) are set. The coordinates are converted to UTM x and y to facilitate metric calculations and maintain the mission within a rectangular boundary. To ensure the mission remains within one or two UTM zones, it is advisable to restrict its scope to a practical distance. The first image is centered at the top left coordinates of the map. Zoom and image resolution will be calculated from the previous section by inputting the camera $FOV$, $AR$, and the drone’s $AGL$. By default the images are not overlapped and the distance between two consecutive images is the width of the image. Overlap can be accounted for, however. In this situation, the distance will be $d^{\prime}=d(1-overlap)$ where $overlap$ is the proportion of shared width/height to the total and has a value between 0 and 1. The mission ends when the bottom right coordinate is captured.
@@ 3 Usage and Results
3.1 SkyAI Dataset
Using SkyAI-Sim, some sample data has been generated and published in this paper available in the "dataset" folder. "Memphis_single_sample" is an example of using the "download_single" program."random_samples" folder includes sample data listed in the "sample_coords.txt". This data is collected using "download_from_list" program. The main dataset includes the raster images of the Memphis Agricenter Area in Figure 7 referred to as Agricenter dataset. The images are collected in the rectangular bound between top left coordinates of $(35.16,-89.90)$ and bottom right coordinates of $(35.115,-89.823)$ using the "download_raster" program. The simulation has been done using a UAV based camera with a $FOV$ of 78.8 degrees and $AR$ of 4:3. The flight takes place at a fixed $AGL$ of 120 m. There are two sets of 1806 images each in "satellite" and "roadmap" modes without any overlap. According to the publish limits from the Google Map side, we could not publish a bigger dataset with this paper, however, a bigger dataset from Memphis area helped with the project described in [14]. This data is placed in "Memphis" folder under two subfolders "satellite_0" and "roadmap_0". This dataset is specifically designed to be used for machine learning tasks. For ease of access, the images are geotagged; a metadata table is also generated named "meta_data.csv". This file contains 6 columns: 1. "img_names" has the name of all images available. 2. "columns" and 3. "rows" consequently refer to the column and row indices of the raster search. 4. "Lat", 5. "Lon" also refer to the central coordinates of each image. 6. "Alt" for this dataset stores the zoom level value for each image. Some corresponding ML processings will be explained as follows.
@@ 3.2 Supplementary Dataset
As previously noted, there are restrictions on publishing data from the API, making the SkyAI datasets themselves, insufficient for machine learning purposes. However, a more extensive dataset can be obtained from publicly available road data. In light of this, we are offering a larger second dataset that covers the city of Memphis. This additional dataset will serve as a valuable resource and can be used for comparison with the data generated by SkyAI Sim in various machine learning tasks.
@@
The bounding box for this data has a top left coordinate of (35.218752, -90.075361), and bottom right (35.064913, -89.730661). The images are taken in an $AGL$ of 300 m with horizontal and vertical $FOV$ of 75 and make a total of 3350 images. This additional dataset will enhance our ability to generate road data and will serve as a key point of comparison with the data from SkyAI Sim. By comparing the two datasets, we can better assess their strengths and limitations, ultimately improving the accuracy of our machine learning models specifically for road extraction tasks. Figure 8 illustrates one sample of this dataset.
@@ 3.3 Usage
To start using the API, you first need a project with a billing account on Google Maps API. When integrating the Google Maps Static API into SkyAI-Sim, it’s crucial to understand the API’s limitations and Google’s policies on content usage  [^9] https://about.google/brand-resource-center/products-and-services/geo-guidelines/#required-attribution/. The code architecture is as follows.
Figure 9: Code Structure.
@@
  • skyai.data: The data package has three main data modules in it:
@@
      SkyAI: This is the main module to load the SkyAI data. It is inherited from VBN module and has its specific attributes and methods to handle GoogleMap static API.
  • skyai.utils: The utility package for managing more generic tasks:
@@
  • skyai.notebooks: There are two main notebooks available in this directory They both take care of processing of the datasets that are currently $d$atacleaner" notebook helps with looking at the entropies of random samples and preparing the histogram and analysis on the data cleaner threshold selection. "preprocess" notebook eases the visualization of applying various preprocessing tools on random sample images from the downloaded data.

### v2

# TerrAInav Sim: An Open-Source Simulation of UAV Aerial Imaging from Satellite Data

@@ Abstract
Capturing real-world aerial images for vision-based navigation (VBN) is challenging due to limited availability and conditions that make it nearly impossible to access all desired images from any location. The complexity increases when multiple locations are involved. State-of-the-art solutions, such as deploying UAVs (unmanned aerial vehicles) for aerial imaging or relying on existing research databases, come with significant limitations. TerrAInav Sim offers a compelling alternative by simulating a UAV to capture bird’s-eye view map-based images at zero yaw with real-world visible-band specifications. This open-source tool allows users to specify the bounding box (top-left and bottom-right) coordinates of any region on a map. Without the need to physically fly a drone, the virtual Python UAV performs a raster search to capture images. Users can define parameters such as the flight altitude, aspect ratio, diagonal field of view of the camera, and the overlap between consecutive images. TerrAInav Sim’s capabilities range from capturing a few low-altitude images for basic applications to generating extensive datasets of entire cities for complex tasks like deep learning. This versatility makes TerrAInav a valuable tool for not only VBN but also other applications, including environmental monitoring, construction, and city management. The open-source nature of the tool also allows for the extension of the raster search to other missions. A dataset of Memphis, TN, has been provided along with this simulator. A supplementary dataset is also provided, which includes data from a 3D world generation package for comparison.
Keywords Vision-Based Navigation  $\cdot$ UAV  $\cdot$ Satellite Image; Aerial Imaging  $\cdot$ Simulation  $\cdot$ Dataset
@@ 1 Introduction
Figure 1: Image captured at coordinates (35.128039, -89.799163) at above ground level altitude ($AGL$) of 126 m. Image captured (a) by a UAV and simulated (b-d) with a field of view ($FOV$) of 78.8 degrees and aspect ratio ($AR$) of 4:3. Figures b-d represent simulation using three different maptypes.
@@ 1.1 Application
Aerial imagery is crucial for diverse applications, including construction, city administration, environmental monitoring, and vision-based navigation. In artificial intelligence (AI) and machine learning (ML), aerial images provide essential data for training models that power autonomous systems, such as drones and self-driving cars, enabling them to navigate by identifying barriers and landmarks [30]. For environmental monitoring, aerial imagery allows for assessing land use changes, tracking deforestation, and observing wildlife habitats, all of which are vital for conservation efforts [15]. In construction, these images assist in project management by offering real-time site overviews, improving resource allocation, and ensuring safety compliance [17]. City management benefits from aerial imagery in urban planning, infrastructure maintenance, and disaster response, providing detailed views of city layouts [7].
Aerial image datasets play a crucial role in machine learning, supporting tasks such as object detection, image classification, and segmentation. These datasets enable the identification of objects like vehicles, buildings, and vegetation, contributing to urban planning, environmental monitoring, and traffic analysis [14]. They also support image classification tasks, such as distinguishing between land cover types like agricultural, residential, and forested areas, which is crucial for resource management [18]. UAV-based remote sensing has been widely applied across domains like agriculture, ecology, and geography. For instance, UAV hyperspectral imaging has been employed for crop yield prediction using machine learning techniques [13], highlighting the potential of UAV-based remote sensing in precision agriculture. While hyperspectral imaging provides high spectral resolution, other approaches, such as RGB or multispectral imaging, offer cost-effective alternatives for similar applications. In semantic segmentation, aerial imagery allows for pixel-level classification, aiding applications such as crop monitoring, disaster response, and infrastructure assessment [23]. Additionally, change detection algorithms use temporal aerial imagery to monitor land feature changes over time, essential for tracking deforestation, urban expansion, and environmental shifts [6]. Furthermore, object tracking is another critical application of aerial imagery, enabling real-time monitoring of moving objects like vehicles and ships in surveillance and security contexts [29].
@@ 1.2.1 Datasets
Acquiring real-world aerial images is challenging due to constraints. The limited availability of images is a primary issue, as it is often impractical to obtain the required aerial views from specific locations. This difficulty is compounded when there is a need to cover multiple locations on Earth. Deploying Unmanned Aerial Vehicles (UAVs) comes with inherent limitations. For instance, flying a UAV to capture aerial images involves logistical complexities, significant time investment, and financial costs. Moreover, UAV operations are subject to regulatory restrictions, such as airspace regulations and flight permissions, as well as environmental factors like weather conditions that can impede flight missions. Numerous aerial image datasets are available, serving applications such as segmentation and object detection. Real-world aerial images captured by drones are the source of several databases. For instance, FloodNet [21, 22], which comprises 2,343 samples of flooded areas for classification, semantic segmentation, and visual question answering (VQA) in flood management. The University1652-Baseline dataset [32, 31] offers satellite-view besides the cross-view (CV) and drone-based images.
While these datasets are valuable for many tasks, their usefulness is constrained by the parameters set by their authors and may not align with the needs of this research. Unlike a real drone mission, these datasets lack the flexibility to select preferred locations and altitudes. They also often do not match particular conditions, such as precise locations and camera specifications. Although useful, they may not cover the scenarios needed for comprehensive testing or algorithm development.
A custom imaging simulator offers a controlled environment to generate synthetic data, customize conditions, and adapt to new models, providing repeatability, safety, and cost-effectiveness while addressing issues like incomplete data and challenging environments that are difficult to replicate with real-world datasets. Moreover, existing UAV-based datasets are often designed for specific applications, such as classification, which limits their adaptability to broader tasks. Even datasets that support multiple applications tend to focus on a narrow set of tasks, restricting their usability in more complex scenarios. Our approach extends these applications by incorporating a wider range of vision-based tasks, including image pattern recognition [10], which, when combined with geotags, can contribute to geolocation algorithms. Additionally, segmentation techniques can be applied for road detection and navigation, while image regression, classification, and image-to-image translation open possibilities for applications such as generating synthetic maps for simulation and game design. For example, UAV imagery has been used for road segmentation to enhance autonomous navigation systems [16], and generative models have been applied to aerial imagery to synthesize realistic landscapes for virtual environments [25]. By expanding the scope of UAV-based image analysis, our dataset aims to bridge these gaps and enable more versatile applications across multiple domains.
1.2.2 Simulators
Flight simulators such as Microsoft Flight Simulator [9], FlightGear Flight Simulator (FGFS) [12], and Geo-FS [27] each come with fundamental limitations that make them unsuitable for imaging research, such as being authentically designed for fixed-wing aircraft flight simulation and limited drone and imaging features that restrict their utility in flight mission imaging research, requiring large-scale raster image collection for machine learning. Microsoft Flight Simulator, while providing a highly detailed and realistic visual experience, is a closed-source platform, preventing any meaningful customization. It is also computationally demanding, requiring powerful hardware, which limits its practicality for high-frequency image capture over large areas. FlightGear, despite being open-source, presents similar challenges. It was not designed for imaging applications and lacks native support for drone operations or systematic raster image capture. Significant customization would be required to adapt it, and even then, it suffers from limited terrain resolution and image realism, especially at low altitudes. "Geo-FS", while lightweight and accessible as a browser-based simulator, imposes further restrictions: it does not support drone simulations or systematic raster image collection, lacks geotagged data outputs, limits the user’s ability to control camera settings or define extensive missions due to its web-based design and reliance on JavaScript, which hinders compatibility with Python for seamless data integration and manipulation.
Among available flight simulators, FGFS is the only candidate that might be capable of providing comparable data, but not in its default state. It lacks built-in support for precise, automated imaging missions, requiring extensive customization and the development of additional tools to extract usable data. Other simulators, however, are entirely unsuitable, as they are designed purely for piloting experiences and lack any framework for aerial image data generation. Their architectures do not accommodate the level of control, automation, and fidelity necessary for machine learning dataset development, making them completely impractical for this purpose.
To overcome some of these limitations, an alternative solution, such as the MSFS Map Enhancement mod [11], replaces Microsoft Flight Simulator’s default maps with Google Maps imagery, offering more realistic and up-to-date terrain representation. This enhancement along with its being open access, makes it a potential asset for imaging research. However, as a third-party modification, Microsoft Flight Simulator is required to function and does not operate as a standalone tool, limiting its flexibility for large-scale data collection. Despite this, it serves as a valuable reference for integrating high-quality map-based imagery for terrAInav Sim.
Developing robust machine learning models for aerial imagery applications requires access to large, diverse datasets that ensure both accuracy and generalization [26]. TerrAInav Sim  [^1] https://github.com/JacobsSensorLab/TerrAInav-Sim addresses these needs by providing synthetic aerial datasets that can be used for data augmentation, a valuable resource when labeled data is scarce or costly to obtain. This lightweight, versatile tool captures geo-tagged, simulated aerial images, emulating a UAV capturing bird’s-eye view map-based images with real-world visible-band specifications. By default, the simulated drone is positioned at a zero-yaw, facing north, with a 90-degree tilt (looking straight down), as shown in Figure 1. While this orientation is limited to a top-down perspective, it aligns well with many applications that rely on bird’s-eye views for analysis. Through TerrAInav Sim, users can simulate diverse missions by specifying coordinates for any global location. Additionally, for users with the means to collect their own drone-view images, TerrAInav extends the capabilities of the University1652 dataset to any region, enabling tasks that integrate both drone-view and satellite-view images for enhanced data collection and analysis.
@@ 2.1.1 Single Image Capture
The process of capturing a map, involves understanding the relationship between the zoom level, image pixel size, and geolocation bounds. The API allows one to specify the zoom level of the map. The zoom level determines the scale of the map and, consequently, the resolution of the captured image. Higher zoom levels provide more detail, while lower zoom levels cover larger areas with less detail. As follows, this section explains how the simulation parameters $AGL,FOV,AR$ are converted to parameters required to capture satellite data including $z$ and $res_{x,y}$  [^2] The functions introduced in ”Single Image Capture” section endnotes are located in the ”src/utils” folder within the ”geo_helper” package..
@@
    The Google Maps JavaScript API [1] getBounds() function provides the latitude and longitude boundaries of the visible map area based on the central coordinates, zoom level, and map size in pixels ($res_{x,y}$, also referred to as the resolution of the map). To simulate a geographic area from a central point, zoom level, and map dimensions, we calculate the bounding box coordinates using Mercator projection principles. The following steps detail this calculation process. 555”calc_bbox_api” function implements bounding box calculation in Python.
@@ 2.1.2 Raster Image Mission
Building on single-image capture, more advanced missions can be defined, such as raster image collection. The main purpose of TerrAInav Sim is to capture raster images of a specified area. In this mission, the objective is to complete a zero-tilt raster image capture, as shown in Figure 2. The capture begins once the bounding box coordinates of the mission’s rectangular region (the map) are set. The coordinates are converted to UTM x and y to facilitate metric calculations and maintain the mission within a rectangular boundary. To ensure the mission remains within one or two UTM zones, it is advisable to restrict its scope to a practical distance. The first image is centered at the top left coordinates of the map. Zoom and image resolution will be calculated from the previous section by inputting the camera $FOV$, $AR$, and the drone’s $AGL$. By default, images do not overlap, and the distance between consecutive captures equals the image width. Overlap can be introduced by adjusting the spacing using $d^{\prime}=d(1-overlap)$ where $overlap$ is the proportion of shared width/height to the total and has a value between 0 and 1. The mission ends when the bottom right coordinate is captured.
@@ 3 Usage and Results
3.1 TerrAInav Dataset
Using TerrAInav-Sim, some sample data has been generated and published in this paper available in the "dataset" folder. "Memphis_single_sample" is an example of using the "download_single" program."random_samples" folder includes sample data listed in the "sample_coords.txt". This data is collected using "download_from_list" program. The main dataset includes the raster images of the Memphis Agricenter Area in Figure 7 referred to as Agricenter dataset. The images are collected in the rectangular bound between top left coordinates of $(35.16,-89.90)$ and bottom right coordinates of $(35.115,-89.823)$ using the "download_raster" program. The simulation has been done using a UAV based camera with a $FOV$ of 78.8 degrees and $AR$ of 4:3. The flight takes place at a fixed $AGL$ of 120 m. There are two sets of 1806 images each in "satellite" and "roadmap" modes without any overlap. A bigger dataset from Memphis area helped with the project described in [10]. This data is placed in "Memphis" folder under two subfolders "satellite_0" and "roadmap_0". This dataset is specifically designed to be used for machine learning tasks. For ease of access, the images are geotagged; a metadata table is also generated named "meta_data.csv". This file contains 6 columns: 1. "img_names" has the name of all images available. 2. "columns" and 3. "rows" consequently refer to the column and row indices of the raster search. 4. "Lat", 5. "Lon" also refer to the central coordinates of each image. 6. "Alt" for this dataset stores the zoom level value for each image. Some corresponding ML processings will be explained as follows.
@@ 3.2 Supplementary Dataset
As previously noted, there are restrictions on publishing data from the API, making the TerrAInav datasets themselves, insufficient for machine learning purposes. However, a more extensive dataset can be obtained from publicly available road data. In light of this, we are offering a larger second dataset that covers the city of Memphis. This additional dataset will serve as a valuable resource and can be used for comparison with the data generated by TerrAInav Sim in various machine learning tasks.
@@
The bounding box for this data has a top left coordinate of (35.218752, -90.075361), and bottom right (35.064913, -89.730661). The images are taken in an $AGL$ of 300 m with horizontal and vertical $FOV$ of 75 and make a total of 3350 images. This additional dataset will enhance our ability to generate road data and will serve as a key point of comparison with the data from TerrAInav Sim. By comparing the two datasets, we can better assess their strengths and limitations, ultimately improving the accuracy of our machine learning models specifically for road extraction tasks. Figure 8 illustrates one sample of this dataset.
@@ 3.3 Usage
  • terrainav.data: The data package has three main data modules in it:
@@
      TerrAInav: This is the main module to load the TerrAInav data. It is inherited from VBN module and has its specific attributes and methods to handle GoogleMap static API.
  • terrainav.utils: The utility package for managing more generic tasks:
@@
  • terrainav.notebooks: There are two main notebooks available in this directory They both take care of processing of the datasets that are currently $d$atacleaner" notebook helps with looking at the entropies of random samples and preparing the histogram and analysis on the data cleaner threshold selection. "preprocess" notebook eases the visualization of applying various preprocessing tools on random sample images from the downloaded data.